OpenStudy (anonymous):

What is the product of eight over fifteen times ten over eleven in simplest form?

OpenStudy (anonymous):

i will give u medal

undeadknight26 (undeadknight26):

8/15 * 10/11?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

darn beat me to answering, can anyone help me with my question?

OpenStudy (anonymous):

16/33

OpenStudy (skullpatrol):

Use the multiplication rule for fractions: Multiply their numerators and multiply their denominators.

OpenStudy (anonymous):

um ok

OpenStudy (skullpatrol):

To simplify divide the numerator and the denominator by 5.

OpenStudy (skullpatrol):

$$\Huge \frac{ac}{bd}=\frac{a}{b}*\frac{c}{d}$$

OpenStudy (anonymous):

i still dont understand

OpenStudy (skullpatrol):

For simplification: $$\Huge \frac{bc}{bd}=\frac{c}{d}$$

OpenStudy (anonymous):

o now i get it. so it would look like this? 8/15 * 10/11 = 8*10 / 15*11 = 80/165= 16 / 33 answer is 16/33

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

ok thnx. ur the best

OpenStudy (skullpatrol):

Thanks for trying to understand :)

OpenStudy (anonymous):

no thank u 4 helping c: